# The Multilingual Mind: A Survey of Multilingual Reasoning in Language Models

### v1

@@ Abstract
While reasoning and multilingual capabilities in Language Models (LMs) have achieved remarkable progress in recent years, their integration into a unified paradigm—multilingual reasoning—is at a nascent stage. Multilingual reasoning requires language models to handle logical reasoning across languages while addressing misalignment, biases, and challenges in low-resource settings. This survey provides the first in-depth review of multilingual reasoning in LMs. In this survey, we provide a systematic overview of existing methods that leverage LMs for multilingual reasoning, specifically outlining the challenges, motivations, and foundational aspects of applying language models to reason across diverse languages. We provide an overview of the standard data resources used for training multilingual reasoning in LMs and the evaluation benchmarks employed to assess their multilingual capabilities. Next, we analyze various state-of-the-art methods and their performance on these benchmarks. Finally, we explore future research opportunities to improve multilingual reasoning in LMs, focusing on enhancing their ability to handle diverse languages and complex reasoning tasks.
@@ 1 Introduction
If we spoke a different language, we would perceive a somewhat different world.         Ludwig Wittgenstein
Figure 1: Overview of different aspects covered in our survey.
Large Language Models (LLMs) have emerged as transformative tools in natural language processing, demonstrating state-of-the-art performance in language generation, translation, and summarization. These models, trained on vast corpora, excel in generating human-like text and understanding diverse linguistic contexts. Despite their success in language generation, LLMs often face significant challenges in addressing underrepresented languages and reasoning.
@@
Reasoning is formally defined as the process of drawing logical conclusions based on available information, enabling individuals and systems to solve problems and make complex decisions. Recent advancements have sought to bridge this gap by enhancing reasoning capabilities in LLMs using Chain-of-Thought (CoT) prompting (Wei et al., 2022), fine-tuning (Lobo et al., 2024), and hybrid modeling (Yao et al., 2024), especially in high-resource languages like English. However, reasoning in multilingual contexts remains a relatively unexplored domain. Existing efforts predominantly focus on a handful of high-resource languages, leaving low-resource and typologically distant languages underrepresented. The lack of robust benchmarks, diverse training corpora, and alignment strategies further impede progress in this vital area.
Multilingual reasoning, which combines logical inference with multilingual capabilities, is essential for creating AI systems that can operate effectively across diverse linguistic and cultural contexts (Shi et al., 2022). Such systems hold immense potential for global applications, from multilingual education to culturally adaptive healthcare, ensuring inclusivity and fairness. The motivation for this survey arises from the urgent need to address these challenges and provide a systematic exploration of methods, resources, and future directions for multilingual reasoning in LLMs. The different aspects covered in this survey are shown in Figure 1 and
Figure 2: Taxonomy tree of current Multilingual Research. The thrusts for improving multilingual reasoning mainly include representation learning, fine-tuning, prompting, and model editing. With the emergence of multilingual LLMs, while initial research focused on naive prompting, recent works propose several alignment, editing, and fine-tuning strategies to improve reasoning in multilingual LLMs.
the contributions are as follows:
@@ 2 Multilingual Reasoning in LLMs
Recent advancements in large language models have improved reasoning in mathematics and logical reasoning. However, extending these abilities across languages introduces several challenges, including consistency, low-resource adaptation, and cultural integration. Below, we describe the preliminaries and key characteristics of multilingual reasoning, focusing on challenges and desiderata for cross-lingual inference.
@@ 2.1 Preliminaries
Large Language Models (LLMs). LLMs are transformer-based (Vaswani, 2017) neural network architectures designed to model the probability of a sequence of tokens. Formally, LLMs are trained to predict the probability of a word (or sub-word token) given the preceding words in a sequence $X=\{x_{1},x_{2},\dots,x_{n}\}$, formalized as:
$P(X)=\prod_{i=1}^{n}P(x_{i}\mid x_{1},x_{2},\dots,x_{i-1}),$
where $P(X)$ is the probability of the entire sequence and $P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$ is the conditional probability of the ith token given the preceding tokens.
Reasoning. One of the key reasons behind the success of LLMs in mathematic and logical tasks is their reasoning capabilities. Formally, reasoning enables LLMs to draw logical conclusions $C$ from premises $P$ using a mapping function: $C=f(P)$. To this end, there are different types of reasoning strategies that an LLM can employ:
a) Deductive Reasoning: It derives specific conclusions from general premises. If a given set of premises $P_{i}$ are true, the conclusion $C$ must be true.
$P_{1},P_{2},...,P_{n}\Rightarrow C,$
b) Inductive Reasoning: Generalizes patterns from specific instances, leading to probabilistic conclusions.
$P_{1},P_{2},...,P_{n}\Rightarrow C_{\text{probabilistic}}$
c) Abductive Reasoning: Infers the most plausible explanation ($H_{\text{best}}$) for given observation $O$.
$O\Rightarrow H_{\text{best}}$
d) Analogical Reasoning: Identifies relationships between domains and transfers knowledge.
$A:B\approx C:D$
e) Commonsense Reasoning: Uses real-world knowledge for intuitive decision-making.
@@ 2.2 Desiderata in Multilingual Reasoning
Here, we describe desiderata that lay the foundation for multilingual reasoning capabilities in LLMs. Multilingual reasoning refers to the capability of models to perform logical inference, problem-solving, and decision-making across multiple languages while maintaining consistency and cultural contextualization. Let $L{=}\{l_{1},l_{2},\dots,l_{m}\}$ represent a set of $m$ languages, and let $P_{l}$ and $C_{l}$ denote the premise and conclusion in a given language $l_{i}$. For a multilingual reasoning model $M$, the task can be defined as: $M(P_{l_{i}})\to C_{l_{i}},\quad\forall l_{i}\in L,$ where $M$ must satisfy the following key desiderata:
1. Consistency: A model should make logically equivalent conclusions across languages for semantically equivalent premises, i.e.,
$C_{l_{i}}\approx C_{l_{j}},\quad\text{if }P_{l_{i}}\equiv P_{l_{j}},\quad\forall l_{i},l_{j}\in L,$
where $\equiv$ indicates semantic equivalence of premises across languages. Consistency ensures that logical conclusions remain invariant of the input language.
2. Adaptability: For languages $l_{k}\in L_{\text{low-resource}}$, the model must generalize effectively using cross-lingual transfer from high-resource languages and perform robust reasoning, i.e.,
$\forall l_{k}\in L_{\text{low-resource}},\quad M(P_{l_{k}})\to C_{l_{k}},$
where transfer is performed from $l_{k}\in L_{\text{high-resource}}$.
3. Cultural Contextualization: Reasoning should consider cultural and contextual differences inherent to each language, i.e., for a context $c_{l_{i}}$ specific to language $l_{i}$, the conclusion $C_{l_{i}}$ should adapt accordingly:
$C_{l_{i}}=f(P_{l_{i}},c_{l_{i}}),\quad\forall l_{i}\in L,$
where $f$ is a mapping function that integrates linguistic reasoning with cultural nuances, critical for tasks like healthcare, policy enforcement, and education, where culturally informed reasoning is required.
4. Cross-Lingual Alignment: The model must align reasoning processes across typologically diverse languages, where typology refers to linguistic differences in syntax, morphology, and structure (e.g., word order variations between English and Japanese). Given the typological variations $T_{l_{i}}$ and $T_{l_{j}}$ for languages $l_{i}$ and $l_{j}$, alignment ensures that reasoning remains consistent and coherent across languages, regardless of their structural differences, i.e.,
$\text{if }P_{l_{i}}\equiv P_{l_{j}},\quad M(P_{l_{i}})\approx M(P_{l_{j}}),\quad\forall l_{i},l_{j}\in L.$
Next, we highlight existing works that propose different training corpus and benchmarks for multilingual reasoning in Sec. 3 and then describe previously proposed techniques to improve multilingual reasoning of LLMs in Sec. 4. The different thrusts of multilingual reasoning research are shown in Figure 2.
@@ 3 Multilingual Reasoning Datasets
The role of multilingual datasets in reasoning tasks is crucial for developing robust and linguistically diverse LLMs. Models trained on monolingual corpora often exhibit language biases (Lyu et al., 2024), limiting their reasoning capabilities across non-English languages. Therefore, multilingual datasets play a vital role in ensuring equitable model performance, particularly in high-stakes domains such as science, law, and healthcare (Hendrycks et al., 2020). Additionally, we need robust benchmarks to evaluate the effectiveness of various LLMs and techniques in handling critical domain-specific reasoning queries across both high-resource and low-resource languages (Xu et al., 2024; Rasiah et al., 2024). Below, we analyze the datasets based on training datasets (Sec. 3.1), and evaluation benchmarks (Sec. 3.2), comprising domains, tasks, and language distribution in current multilingual reasoning datasets.
@@ 3.1 Training Corpus
The best strategy to equip a language model with a specific type of reasoning is to train the model on it. However, the training objective differs based on the use case, domain, and the language in which the model needs to be adapted. For example, to perform some form of mathematical reasoning (e.g., GSM8K (Cobbe et al., 2021) and OpenMathQA (Amini et al., 2019) are large datasets that are used to improve mathematical reasoning) in a particular language, it needs to be trained with mathematical reasoning datasets, which will differ if we want to adapt the model for legal reasoning.
Figure 3: Language distribution across training corpora and benchmarks for multilingual reasoning. The y-axis denotes the number of training corpora/benchmark datasets that include a given language (x-axis). We observe a long-tail distribution denoting that current datasets predominantly cover languages like Chinese, English, French, and German, highlighting the need for benchmarks that represent long-tail languages.
While most training corpora are predominantly based on mathematical reasoning, XCSQA (Zhu et al., 2024) and MultiNLI (Williams et al., 2017) are used for enhancing logical and coding reasoning and sPhinX (Ahuja et al., 2024) is developed to translate instruction-response pairs into 50 languages for fine-tuning. In addition, there are cases where translation datasets like OPUS (Tiedemann, 2012), FLORES-200 (Goyal et al., 2022), and LegoMT (Yuan et al., 2022) are used to map the multilingual representation into the representation space of LLMs. Further, Ponti et al. (2020) introduced XCOPA, showing that multilingual pre-training and zero-shot fine-tuning under-perform compared to translation-based transfer. We argue that, moving forward, selecting the appropriate dataset and training methodology is crucial for optimizing a model’s performance in specialized reasoning tasks.
@@ 3.2 Evaluation Benchmark
Evaluation benchmarks are fundamental to advancing the field of multilingual reasoning as they provide a consistent and systematic framework to assess the performance of models across diverse reasoning tasks, such as logical inference, mathematical problem-solving, and cross-lingual understanding. Each reasoning task and domain presents unique challenges, making it crucial to have tailored benchmarks that reflect the specific requirements and complexities of those tasks. Below, we analyze the evaluation benchmarks on three key aspects namely domain (Fig. 4), task (Fig. 5), and languages (Fig. 3) in detail.
@@ 3.2.1 Domains and Tasks Covered
Figure 4: Distribution of multilingual reasoning datasets. We observe that datasets predominantly comprise logical, commonsense, and math reasoning datasets, and the community needs more benchmarks to include temporal, compositional, and tabular reasoning.
Multilingual reasoning in language models spans multiple domains, each with its own set of complexities and requirements. Understanding these differences is essential for developing language models that can effectively adapt to various applications. For instance, Cobbe et al. (2021) highlighted that mathematical reasoning requires structured multi-step logic and datasets like GSM8K. While Ponti et al. (2020) showed that causal reasoning in XCOPA relies on cross-lingual consistency and commonsense inference, Tiedemann (2012) noted that multilingual reasoning introduces typological challenges. These studies emphasize the need for tailored approaches to address the specific demands of each task and domain. Hence, it is crucial to build reliable and robust benchmarks for all domains and tasks to develop more robust techniques tailored to handle the complexity of a particular domain and task. Figures 4-5 show the distribution of datasets across various domains and tasks, highlighting the need to develop more comprehensive benchmarks across multiple domains. Currently, tasks such as math, legal, and commonsense reasoning dominate multilingual evaluation benchmarks, collectively accounting for 54% of the total (see Fig. 5). In contrast, domains like science, visual reasoning, tabular reasoning, temporal reasoning, and ethics are underrepresented, covering only 35%. Furthermore, while commonsense reasoning makes a significant contribution to existing works, other reasoning tasks lag. Notably, crucial domains such as finance and healthcare still lack dedicated evaluation benchmarks for multilingual reasoning, highlighting a significant gap in the field.
@@ 3.2.2 Languages Covered
Figure 6: Taxonomy of Multilingual Reasoning Methods. A categorization of approaches for enhancing multilingual reasoning in language models, covering (A) Representation Alignment, (B) Finetuning, (C) Prompting, and (D) Model Editing.
Comprehensive language coverage is vital for multilingual reasoning, ensuring inclusivity and balanced performance across high-resource and low-resource linguistic communities. Based on languages, current benchmarks can be primarily classified into human languages and coding languages. Benchmarks like XNLI (Conneau et al., 2018), mCSQA (Sakai et al., 2024), and MARC (Keung et al., 2020) predominantly focus on high-resource languages like English, Chinese, French, and Spanish. While some efforts include low-resource languages like Swahili (XCOPA (Ponti et al., 2020)), Haitian (M4U (Wang et al., 2024)), and Nepali (mMMLU (Hendrycks et al., 2020)), their representation remains minimal and research in these languages remain at a nascent stage. Typologically distant and underrepresented languages, such as Kannada, Gujarati (xSTREET (Li et al., 2024a)), and Quechua, are rarely included, further widening linguistic inequities. Datasets like FLORES-200 attempt to balance high- and low-resource languages but fail to achieve comprehensive coverage. We note that to ensure LLMs perform effectively across diverse linguistic and cultural contexts, it is critical to include a broader range of low-resource and endangered languages (Goyal et al., 2022; Amini et al., 2019). The complete distribution of human languages across benchmarks is shown in Figure 3. Finally, only three benchmarks (BBH-Hard, CRUXEval, and TCC) incorporate coding languages across multiple languages.

### v2

# A Survey of Multilingual Reasoning in Language Models

@@ Abstract
While reasoning and multilingual capabilities in Language Models (LMs) have achieved remarkable progress in recent years, their integration into a unified paradigm—multilingual reasoning—is at a nascent stage. Multilingual reasoning requires language models to handle logical reasoning across languages while addressing misalignment, biases, and challenges in low-resource settings. This survey provides the first in-depth review of multilingual reasoning in LMs. In this survey, we provide a systematic overview of existing methods that leverage LMs for multilingual reasoning, specifically outlining the challenges, motivations, and foundational aspects of applying language models to reason across diverse languages. We provide an overview of the standard data resources used for training multilingual reasoning in LMs and the evaluation benchmarks employed to assess their multilingual capabilities. Next, we analyze various state-of-the-art methods and their performance on these benchmarks. Finally, we explore future research opportunities to improve multilingual reasoning in LMs, focusing on enhancing their ability to handle diverse languages and complex reasoning tasks. Rapid growth of evolving developments in this field can be actively tracked on our project page: https://github.com/AkashGhosh/Survey-of-Multilingual-Reasoning-in-Language-Models
A Survey of Multilingual Reasoning in Language Models
Akash Ghosh1thanks: Equal contribution. Work done while interning at Aikyam Lab (UVA). Contact author: akash_2321cs19@iitp.ac.in   Debayan Dutta [^1] footnotemark: 1   Sriparna Saha1   Chirag Agarwal2 1Indian Institute of Technology Patna, India 2University of Virginia, USA
@@ 1 Introduction
If we spoke a different language, we would perceive a somewhat different world.
Ludwig Wittgenstein
Large Language Models (LLMs) (Vaswani, 2017) have emerged as transformative tools in natural language processing, demonstrating state-of-the-art performance in language generation, translation, and summarizationJain et al. (2022); Ghosh et al. (2024a, d, b, 2025); Ghosal et al. (2025). These models, trained on vast corpora, excel in generating human-like text and understanding diverse linguistic contexts. Despite their success in language generation, LLMs often face significant challenges in addressing underrepresented languages and reasoning.
@@
Reasoning is formally defined as the process of drawing logical conclusions, enabling individuals and systems to solve problems and make complex decisions. Recent advancements have sought to enhance the reasoning capabilities of LLMs using Chain-of-Thought (CoT) (Wei et al., 2022), fine-tuning (Lobo et al., 2024), and hybrid modeling (Yao et al., 2024), especially in high-resource languages like English. However, reasoning in multilingual contexts remains a relatively unexplored domain, where existing efforts predominantly focus on a handful of high-resource languages, leaving low-resource and typologically distant languages underrepresented. The lack of robust benchmarks, diverse training corpora, and alignment strategies further impede progress in this vital area.
Multilingual reasoning, which combines logical inference with multilingual capabilities, is essential for creating AI systems that effectively operate across diverse linguistic and cultural contexts (Shi et al., 2022). Such systems hold immense potential for global applications, from multilingual education to culturally adaptive healthcare, ensuring inclusivity and fairness. The motivation for this survey arises from the urgent need to address these challenges and provide a systematic exploration of methods, resources, and future directions for multilingual reasoning in LLMs. The key contributions of our work are:
Figure 1: Taxonomy tree of current Multilingual Reasoning Research. The thrusts for improving multilingual reasoning mainly include representation learning, fine-tuning, prompting, and model editing. With the emergence of multilingual LLMs, while initial research focused on naive prompting, recent works propose several alignment, editing, and fine-tuning strategies to improve reasoning in multilingual LLMs.
@@ 2 Multilingual Reasoning in LLMs
Recent advancements in LLMs have improved their reasoning capabilities. However, extending them across languages introduces several challenges, including consistency, low-resource adaptation, and cultural integration. Below, we describe the preliminaries and key characteristics of multilingual reasoning, focusing on challenges and desiderata for cross-lingual inference.
@@ 2.1 Preliminaries
Large Language Models (LLMs). LLMs are transformer-based neural network architectures designed to model the probability of a sequence of tokens. Formally, LLMs are trained to predict the likelihood of a word (or sub-word token) given the preceding words in a sequence $X=\{x_{1},\dots,x_{n}\}$, i.e., $P(X)=\prod_{i=1}^{n}P(x_{i}\mid x_{1},\dots,x_{i-1}),$ where $P(X)$ is the probability of the entire sequence and $P(x_{i}|x_{1},\dots,x_{i-1})$ is the conditional probability of the ith token given the preceding tokens.
Reasoning. One of the key reasons behind the success of LLMs in mathematical and logical tasks is their reasoning capabilities. Formally, reasoning enables LLMs to draw logical conclusions $C$ from premises $P$ using a mapping function: $C=f(P)$. To this end, there are different types of reasoning strategies that an LLM can employ:
a) Deductive Reasoning: Derives logically certain conclusions from general premises. If the premises $P_{i}$ are true, the conclusion $C$ must also be true, i.e., $P_{1},P_{2},\dots,P_{n}\;\Rightarrow\;C.$
b) Inductive Reasoning: Infers general rules or patterns from specific observations, leading to conclusions that are likely but not guaranteed, i.e., $P_{1},P_{2},\dots,P_{n}\;\Rightarrow\;C_{\text{probabilistic}}.$
c) Abductive Reasoning: Infers the most plausible hypothesis ($H_{\text{best}}$) that explains an observation $O$, though the inference is not guaranteed to be correct, i.e., $O\;\Rightarrow\;H_{\text{best}}.$
d) Analogical Reasoning: Transfers knowledge by identifying relational similarities between domains, i.e., $A:B\;\approx\;C:D.$
e) Commonsense Reasoning: Draws on background knowledge of everyday situations to make intuitive, contextually appropriate inferences.
@@ 2.2 Desiderata in Multilingual Reasoning
Here, we describe desiderata that lay the foundation for multilingual reasoning in LLMs. Let $L{=}\{l_{1},l_{2},\dots,l_{m}\}$ represent a set of $m$ languages, and let $P_{l}$ and $C_{l}$ denote the premise and conclusion in a given language $l_{i}$. For a multilingual reasoning model $M$, the task can be defined as: $M(P_{l_{i}})\to C_{l_{i}},\quad\forall l_{i}\in L,$ where $M$ must satisfy the following key desiderata:
1. Consistency: A model should make logically equivalent conclusions across languages for semantically equivalent premises, i.e., $C_{l_{i}}\approx C_{l_{j}},\quad\text{if }P_{l_{i}}\equiv P_{l_{j}},\quad\forall l_{i},l_{j}\in L,$ where $\equiv$ indicates semantic equivalence of premises across languages. Consistency ensures that logical conclusions remain invariant of the input language.
2. Adaptability: For languages $l_{k}\in L_{\text{low-resource}}$, the model must generalize effectively using cross-lingual transfer from high-resource languages and perform robust reasoning, i.e., $\forall l_{k}\in L_{\text{low-resource}},\quad M(P_{l_{k}})\to C_{l_{k}}.$
3. Cultural Contextualization: Reasoning should consider cultural and contextual differences inherent to each language, i.e., for a context $c_{l_{i}}$ specific to language $l_{i}$, the conclusion $C_{l_{i}}$ should adapt accordingly: $C_{l_{i}}=f(P_{l_{i}},c_{l_{i}}),\quad\forall l_{i}\in L,$ where $f$ is a mapping function that integrates linguistic reasoning with cultural nuances.
4. Cross-Lingual Alignment: The model must align reasoning processes across typologically diverse languages, where typology refers to linguistic differences in syntax, morphology, and structure (e.g., word order variations between English and Japanese). Given the typological variations $T_{l_{i}}$ and $T_{l_{j}}$ for languages $l_{i}$ and $l_{j}$, alignment ensures that reasoning remains consistent and coherent across languages, i.e., $\text{if }P_{l_{i}}\equiv P_{l_{j}},\quad M(P_{l_{i}})\approx M(P_{l_{j}}),\quad\forall l_{i},l_{j}\in L.$
Next, we highlight existing works that propose different training corpora and benchmarks for multilingual reasoning in Sec. 3 and then describe previously proposed techniques to improve the multilingual reasoning of LLMs in Sec. 4.
@@ 3 Multilingual Reasoning Datasets
Models trained on english corpora exhibit language biases (Lyu et al., 2024), limiting their reasoning capability on non-English languages. Training an LM to solve math problems across languages requires multilingual understanding and mathematical reasoning (Son et al., 2024). Hence, multilingual datasets and benchmarks play a key role in training multilingual LMs and evaluating the effectiveness of various LMs and techniques in handling domain-specific reasoning queries across low- and high-resource languages (Xu et al., 2024; Rasiah et al., 2024; Xue et al., 2024). Below, we detail training datasets (Sec. 3.1) and benchmarks (Sec. 3.2), comprising domains, tasks, and language distribution in current multilingual reasoning datasets.
@@ 3.1 Training Corpus
The best strategy to equip an LM with a specific type of reasoning is to train the model on it. However, the training objective differs based on the use case, domain, and language in which the model needs to be adapted. For example, to perform mathematical reasoning (Cobbe et al., 2021; Amini et al., 2019) in a particular language, it needs to be trained with mathematical reasoning datasets, which will differ if we want to adapt the model for legal reasoning.
Figure 2: Language distribution across training corpora and benchmarks for multilingual reasoning. The y-axis denotes the number of training corpora/benchmark datasets that include a given language (x-axis). We observe a long-tail distribution, denoting that current datasets predominantly cover languages like Chinese, English, French, and German, highlighting the need for benchmarks that represent long-tail languages.
While most training corpora are predominantly based on mathematical reasoning, XCSQA (Zhu et al., 2024b) and MultiNLI (Williams et al., 2017) are used for enhancing logical and coding reasoning, and sPhinX (Ahuja et al., 2024) is developed to translate instruction-response pairs into 50 languages for fine-tuning. In addition, there are cases where translation datasets like OPUS (Tiedemann, 2012), FLORES-200 (Goyal et al., 2022), and LegoMT (Yuan et al., 2022) are used to map the multilingual representation into the LM’s representation space. Further, Ponti et al. (2020) introduced XCOPA to show that multilingual pre-training and zero-shot fine-tuning underperform compared to translation-based transfer. We argue that, moving forward, selecting the appropriate dataset and training methodology is crucial for optimizing a model’s performance in specialized reasoning tasks.
@@ 3.2 Evaluation Benchmark
Benchmarks are key to advancing the field of multilingual reasoning as they provide a systematic framework to assess the performance of models across diverse reasoning tasks. Each reasoning task and domain presents unique challenges, making it crucial to have tailored benchmarks that reflect specific requirements and complexities of those tasks. Below, we analyze the evaluation benchmarks on three key aspects, namely languages (Fig. 2), domain (Fig. 3), and task (Fig. 4).
@@ 3.2.1 Domains and Tasks Covered
Figure 3: Distribution of multilingual reasoning datasets. We find that datasets predominantly comprise logical, commonsense, and math reasoning, and the community needs benchmarks to include compositional and tabular reasoning.
@@
Multilingual reasoning in LMs spans multiple domains, each with its complexities and requirements, and understanding these differences is essential for developing LMs that can effectively adapt to various applications. For instance, Cobbe et al. (2021) highlighted that mathematical reasoning requires structured multi-step logic and datasets. While Ponti et al. (2020) showed that causal reasoning in XCOPA relies on cross-lingual consistency and commonsense inference, Östling and Tiedemann (2016) noted that multilingual reasoning introduces typological challenges. These studies emphasize the need for tailored approaches to address the specific demands of each task and domain. Hence, it is crucial to build reliable and robust benchmarks for developing more robust techniques tailored to handle the complexity of a particular domain and task. Figs. 3-4 show the distribution of datasets across various domains and tasks, highlighting the need to develop more comprehensive benchmarks across multiple domains. Currently, tasks such as math, legal, and commonsense reasoning dominate multilingual benchmarks, collectively accounting for 54% of the total (Fig. 4). In contrast, domains like science, ethics, and visual, tabular, and temporal reasoning are underrepresented, covering only 35%. Notably, crucial domains such as finance and healthcare still lack dedicated evaluation benchmarks for multilingual reasoning, highlighting a significant gap in the field.
@@ 3.2.2 Languages Covered
Figure 5: Taxonomy of Multilingual Reasoning Methods. A taxonomy of approaches for enhancing multilingual reasoning in models, covering (A) Representation Alignment, (B) Finetuning, (C) Prompting, and (D) Model Editing.
Comprehensive language coverage is vital for multilingual reasoning, ensuring inclusivity and balanced performance across low- and high-resource linguistic communities. Based on languages, current benchmarks can be primarily classified into human and coding languages. Benchmarks like XNLI (Conneau et al., 2018), mCSQA (Sakai et al., 2024), and m-ARC (Lai et al., 2023) predominantly focus on high-resource languages like English, Chinese, French, and Spanish. While some efforts include low-resource languages like Swahili (XCOPA (Ponti et al., 2020)), Haitian (M4U (Wang et al., 2024)), and Nepali (mMMLU (Hendrycks et al., 2020)), their representation remains minimal and research in these languages remains at a nascent stage. Typologically distant and underrepresented languages, such as Kannada, Gujarati (xSTREET (Li et al., 2024a)), and Quechua, are rarely included, further widening linguistic inequalities. Datasets like FLORES-200 attempt to balance low- and high-resource languages but fail to achieve comprehensive coverage. To ensure effective LLM performance across diverse linguistic and cultural contexts, it is critical to include a broader range of low-resource and endangered languages (Goyal et al., 2022; Amini et al., 2019) (see the complete distribution of human languages across benchmarks in Fig. 2). Finally, only four benchmarks (Luo et al., 2024; Xu et al., 2024; Zhang et al., 2024b; Li et al., 2024a) incorporate coding languages across multiple languages.